OpenStudy (anonymous):

Which best describes the solution for this system? 2x+2y=4 Y=-x+2 A. There is one unique solution B. There is not a solution C. There are an infinite number of solutions D. The solution is the origin

OpenStudy (homeworksucks):

2x+2y=4 y=-x+2 substituting y=-x+2 into the equation, we get 2x+2(-x+2)=4 expanding that 2x+(-2x)+4=4 2x-2x+4=4 WHAAAAA? No solution.